Find the equation of the line that passes through (-7,2) and (-4,5) in the form Ax+By=C.
Answer by stanbon(75887)

Find the equation of the line that passes through
(-7,2) and (-4,5) in the form Ax+By=C.
-------
slope = (5-2)/(-4--7) = 3/3 = 1
------
Form:: y = mx + b
------
Solve for "b":
5 = 1*-4 + b
b = 9
------
Equation:
y = x + 9
-----
In the form Ax + By = c
x - y = -9
==============================
Cheers,
Stan H.
========================